OpenStudy (anonymous):

Find the Cartesian inequation for the region represented by |z+8|<|z-6| ... Could anyone please perhaps help me with this??

OpenStudy (anonymous):

This was my working, and it was wrong.... Let z=x+iy :. z+8=(x+8)+iy z-6=(x-6)+iy and so |z+8|=sqrt[(x+8)^2 + y^2] |z-6|=sqrt[(x-6)^2 +y^2] goes to (x+8)^2 + y^2 < (x-6)^2 + y^2 (x+8)(x+8) < (x-6)(x-6) which simplifies to 18x < -28 x < -14/9 What have I done incorrectly?

OpenStudy (dumbcow):

(x+8)^2 < (x-6)^2 x^2 +16x +64 < x^2 -12x +36 28x < -28 x < -1

OpenStudy (anonymous):

Oh wow thats annoying hahaha I've been looking at this for ages! Cheers

OpenStudy (dumbcow):

:)